OpenStudy (anonymous):

Choose the missing number to make equivalent fractions. 5/6 20/[]

OpenStudy (anonymous):

Cross multiply

OpenStudy (anonymous):

Cross multiply: \(\dfrac{5}{6} = \dfrac{20}{x}\) What's 5 * x? What's 6 * 20?

OpenStudy (anonymous):

6*20=120

OpenStudy (anonymous):

how do i times x with 5

OpenStudy (anonymous):

5x^2?

OpenStudy (anonymous):

Im confused on how to do that

OpenStudy (anonymous):

No 5 * x is just 5x

OpenStudy (anonymous):

So we have: 5x = 120 Now divide 120 / 5

OpenStudy (anonymous):

24

OpenStudy (anonymous):

Yes that's our answer